Solve:
log(x+1) + log(x-1) = log 48

Answer by lwsshak3(6761)

Solve:
log(x+1) + log(x-1) = log 48
..
log(x+1) + log(x-1)-log 48 = 0
log(x+1)(x-1)/48=0
change to exponential form
10^0=(x+1)(x-1)/48
1=(x+1)(x-1)/48
x^2-1=48
x=sqrt(49)
x=7
x=-7 (reject, x>-1)